OpenStudy (anonymous):

If the function "f(x)=cos(Ax)" has period 3/4, give value of A. Please explain!

OpenStudy (anonymous):

normally the period of cos(x) is 2pi, but when you place a number in front of the x, it can make the period shorter or longer. For example, if I look at the period of cos(2x), its only pi, since multiplying by two makes the x increase twice as fast. In general, if I put cos(Ax), the period would be:\[\frac{2\pi}{A}\]

OpenStudy (anonymous):

so as joe made it clear, your job is to solve \(\frac{2\pi}{A}=\frac{3}{4}\) for \(A\)

OpenStudy (anonymous):

thanks!